# Off-Path TCP Hijacking in Wi-Fi Networks: A Packet-Size Side Channel Attack

### v4

@@ 5.2 TCP Manipulation Attack
Experimental Setup. This attack involves three hosts: a web server, a victim client (our laptop), and an off-path attacker. We use a real financial website [^5] For ethical considerations, we anonymize this financial website in the paper. Moreover, our attack do not affect the website, since we only manipulate the web cache of the client side, i.e., our controlled laptop. as the web server. This website employs HTTP to deliver real-time Bitcoin price to the client in JSON format at 5-second intervals. Before launching the attack, the attacker can explore this website to familiarize themselves with the JSON data structure. Consequently, the attacker can discern specific data within the packet, as depicted in Figure 9(a), enabling manipulation of the victim’s web page. The victim client browses financial information on the web page via Wi-Fi. Both the attacker and the victim client are connected to the same Wi-Fi network. The off-path attacker attempts to detect and hijack the TCP connection between the victim client and the server. The server maintains a single long-lived TCP connection [^6] As recommended in RFC 2616 [39], the client typically does not maintain multiple long-lived TCP connections with the server simultaneously. with the client to transmit real-time financial data. Note that the modern browser may open multiple concurrent TCP connections along with the long TCP connection to speed up the page loading. These concurrent TCP connections are short-lived and have minimal influence on inferring the target long-lived TCP connection. Even if the server maintains multiple long-lived TCP connections with the client in some cases, the attacker can infer all the TCP connections and inject malicious data.
Attack Procedure. The web connection hijacking attack consists of five steps: (i) The attacker determines the MAC address and IP address of the victim client in the WLAN. (ii) By exploiting the encrypted frames, the attacker detects the client’s port number to obtain the TCP 4-tuple information. (iii) The attacker infers the exact sequence number and (iv) gets an acceptable acknowledgment number. (v) The attacker impersonates the server and injects forged TCP packets with the inferred values into the victim client. Finally, the client will accept the forged TCP packets, which subsequently update the financial information on the web page.11
Figure 8: Time/Bandwidth overheads of web manipulation.
@@
Results Evaluation. Figure 8 illustrates the time cost and bandwidth consumption during the attack. It takes an average of 10.1 seconds to identify the client port number and 8.3 seconds to find the exact sequence number. Time cost required to find an acceptable acknowledgment number takes 9.6 seconds. The average duration of the entire attack is 28 seconds, with an average bandwidth cost of 46.3 KB/s. In this case, the attacker needs to infer an acceptable acknowledgment number, hence the success rate of this attack is lower than the TCP DoS attack but still exceeds 70%. After obtaining all the necessary information, the attacker sends forged TCP packets to the victim client and manipulates sensitive data on the web page. Figure 9(b) shows a snapshot of the manipulated web page and where attacker alters the Bitcoin price.
Table 3: Details of 30 tested wireless routers.
@@
6 Real-world Attacks
To assess the impact of our attack, we conduct an extensive investigation on 30 popular wireless routers and 80 real-world Wi-Fi networks. We analyze 30 popular wireless routers and find that all the evaluated routers cannot protect the victim from our attack. Besides, we conduct SSH DoS attack and web hijack attack on the victim (i.e., our device) in the real-world Wi-Fi networks following the experimental setup and procedure in Section 5. The results reveal that our attack is successful [^7] We conduct 10 iterations of the SSH DoS attack and the web hijack attack on the victim. In this context, the “successful” means that these two attacks can be successfully executed at least once in the real-world Wi-Fi network. against 75 (93.75%) of the 80 assessed Wi-Fi networks.
6.1 Analysis of AP Routers
@@
The early WEP used RC4 algorithm for data encryption, whereas WPA2 replaced them with the AES-CCMP algorithm. In the latest WPA3 standard, AES-GCMP is proposed to be used as the encryption method for WPA3 Enterprise mode. However, none of these security mechanisms can prevent the encrypted frame size from leaking upper layer information. Based on our empirical findings, we confirm that all 30 evaluated wireless routers could not protect the supplicant from our attacks.
Table 3 shows detailed information on 30 tested wireless routers in our investigation. Take the first line for example, the evaluated router “Mi 4C” manufactured by Xiaomi belongs to the older Wi-Fi generation (Wi-Fi 4) and lacks support for IPv6 and WPA3. As outlined in the product description, the “Mi 4C” device offers support for various security features. These include a built-in firewall that allows administrators to define packet forwarding rules, a flood defense mechanism that restricts malicious flood traffic to prevent DoS attacks, and MAC address filtering, which enables network access authorization based on hardware addresses. In our investigation, all tested routers claim to support different security mechanisms to prevent various attacks. However, our study demonstrates that the existing security mechanisms are inadequate against our attack.
Table 4: Experimental results in 30 real-world Wi-Fi networks.
@@
6.2 Real-world Wi-Fi Networks Evaluation
The Wi-Fi scenarios we tested cover a wide range of public settings, including coffee shops, restaurants, hotels, cinemas, and bookstores. The experimental results illustrate that over 93% (i.e., 75 out of 80) of the evaluated Wi-Fi networks are vulnerable to our attack. Next, we elaborate on the evaluation results.
@@ 6.2 Real-world Wi-Fi Networks Evaluation
The 802.11n/ac standards are predominantly utilized (73.75%) in real-world Wi-Fi networks. This indicates that these Wi-Fi networks support two frequency bands (i.e., 2.4 GHz and 5 GHz), with their physical layer models based on the 802.11n and 802.11ac standards, respectively. There is only 17.5% (14 out of 80) of the evaluated Wi-Fi networks support 802.11ax. This is consistent with our expectations, as the 802.11ax standard was defined in 2019 and would require more time for widespread deployment. Furthermore, the simultaneous utilization of two channels in Wi-Fi networks is the most prevalent case (76.25%) due to the widespread support and default configuration of dual-channel capabilities in wireless routers. In certain scenarios, such as office buildings, Wi-Fi networks employ multiple wireless channels to enhance network performance. In our study, we identify 12 Wi-Fi networks that utilize multiple wireless channels. At first glance, the usage of multiple wireless channels might appear as a minor hurdle to our attack, as the attacker needs to perform additional channel scanning to determine the specific channel employed by the victim. Indeed, the attacker can enhance the success rate by employing a channel “eviction” strategy, which will be explained in detail in Section 7. Additionally, we encounter several Wi-Fi networks that operate on a single channel. When questioned, network administrators cited security considerations as the rationale behind this choice, although they did not provide any further specifics.
@@
Out of all the evaluated networks, 16 (20%) of them are open and do not encrypt the supplicant’s data frames. This means that an attacker can potentially access and view the contents of all supplicant’s frames transmitted on these networks, representing a significant breach of supplicant privacy. The remaining 64 (80%) Wi-Fi networks utilize WPA2/WPA3 to encrypt the supplicant’s wireless frame. It is worth noting that out of the 80 Wi-Fi networks evaluated, 11 (13.75%) of them have AP isolation enabled. In these Wi-Fi networks, the attacker obtains the victim’s MAC and IP address by leveraging the DHCP mechanism and injects malicious packets into the victim through the AP’s external port. However, our attack encounters failure in five Wi-Fi networks. Among them, one network is equipped with reverse path authentication [51, 58], preventing the attacker from sending packets from the WLAN to the AP’s external port. In the remaining four Wi-Fi networks, the attacker cannot obtain the AP’s external port as the gateway does not respond to ICMP ping messages.
@@
We elaborate more on the experimental results of 30 encrypted Wi-Fi networks in Table 4. We take the first row of Table 4 as an example to analyze the results. In our study, the SSID “Bookstore 1” indicates a Wi-Fi network that is accessible in a bookstore. It is common practice to set the Wi-Fi SSID as the organization name, which may expose the organization’s identity. Therefore, to protect anonymity, we have anonymized the Wi-Fi SSID in this paper. This bookstore’s Wi-Fi network only supports IPv4 and does not have AP isolation enabled, while its AP is produced by ADSLR. This AP provides two access channels (i.e., 6 and 161) and employs the 802.11n and 802.11ac standards. The TCP connection of the victim supplicant can be hijacked using the attack presented in Section 4. The success rates for conducting SSH DoS and web hijacking on the victim supplicant are 70% and 60%, respectively.
@@
Within the evaluated vulnerable Wi-Fi networks, we have observed a range of success rates for our attack, varying from 30% to 80%. The principal factor influencing this variance is the heterogeneous wireless environments (e.g., different wireless interference and channel contention) encountered in real-world Wi-Fi networks, leading to varying capabilities for attacker to capture the victim’s Wi-Fi frames. Factors such as wireless interference (e.g., from microwave ovens and Bluetooth devices) and channel contention can hinder the attacker’s ability to capture the victim’s Wi-Fi frames, resulting in the failure of attacks. For instance, as illustrated in Table 4, Coffee Shop 1, situated on a university campus, experiences less wireless interference and channel contention compared to Restaurant 3, located within a large shopping mall. Consequently, the success rate of attacks in the latter Wi-Fi network is lower due to the elevated interference and contention in that environment. We will conduct a more comprehensive analysis of the factors influencing the success of our attack in Section 7.
7 Discussion
@@
Channel Contention. APs and supplicants based on the 802.11 standards use Carrier Sense Multiple Access with Collision Avoidance (CSMA/CA) to compete equally for the occupation of the wireless channel. Before transmitting frames, wireless channel listening is conducted to ensure that the channel is not occupied. Frames are transmitted only after verifying the channel’s availability. Due to channel contention, there may be an uncertain delay or even frame dropping in the victim’s responses to the probe packets. This uncertain response delay or frame dropping is the primary reason for the fluctuating success rate of our attack because the attacker needs to analyze the victim’s encrypted frames within a time slice after the probe packets are sent. To mitigate channel contention, we propose a channel “eviction” strategy. The attacker can evict other supplicants from the channel used by the victim. Specifically, the attacker impersonates the AP and sends decertification frames to the supplicant, causing it to detach from the current channel of the AP [^9] In the Wi-Fi network with MFP (Management Frame Protection) enabled, attackers can exploit implementation vulnerabilities to force the supplicant to detach from the current channel [49, 48].. The supplicant will attempt to reconnect to the Wi-Fi network, but after encountering several disconnections, it will switch to another channel. This strategy requires the Wi-Fi network to support multiple access channels. Fortunately, most Wi-Fi networks provide more than one access channel, as shown in Section 6.2. Note that the channel switching (i.e., our “eviction” strategy to cause other supplicants to detach from the current channel) is transparent to the users. The only impact is that the user may experience a brief (a few seconds) network jitter during the channel switching.
@@ 9 Related Work
In TCP connection hijacking attacks, side channels serve as potent tools for attackers. The IPID, in particular, has been a frequent target for exploitation. For example, Jeffrey et al. utilized per-destination IPID counters to estimate the number of packets transmitted between two machines and even detect the presence of a TCP connection [35]. Similarly, Alexander et al. used the IPID of triggered RST packets to identify the existence of the victim’s TCP connection [7]. In a recent instance, Feng et al. manipulated the IPID assignment using ICMP to hijack the victim’s TCP connection [22]. Besides IPID, the challenge ACK mechanism is another side channel exploited by off-path attackers. Cao et al., for example, utilized the global rate limit of challenge ACK to infer and hijack TCP connections [14, 15]. Moreover, a timing side channel has been found in half-duplex Wi-Fi technology, which can be exploited by off-path attackers to inject data into the victim’s TCP connection [16]. Tolley et al. recently proposed a blind in/on-path attack in VPNs, aiming to infer the existence of, interfere with, or inject data into TCP connections forwarded through encrypted VPN tunnels [61]. Different from previous research, our work reveals a new side channel, i.e., information leakage due to the encrypted frame size in Wi-Fi networks. This side channel can be exploited by a pure off-path attacker to hijack victim’s TCP connections.
10 Conclusion
@@
In this paper, we present a new off-path TCP hijacking attack that takes advantage of the encrypted frame size in Wi-Fi networks to detect and hijack TCP connections belonging to a victim supplicant. This side channel (i.e., the observable frame size) vulnerability is an inherent flaw in Wi-Fi networks, specifically the 802.11 standards. To execute our attack, attackers initially scan the WLAN to identify active victim supplicant, then analyze the victim’s encrypted frame size to infer the 4-tuple, exact sequence number, and acceptable acknowledgment number of the victim’s TCP connection. Specifically, our attacker can hijack the victim’s TCP connection within 28 seconds. We carry out our attack in typical Wi-Fi scenarios, and our evaluation demonstrates that this new off-path TCP hijacking attack can result in significant damage to upper-layer applications, such as SSH DoS and the injection of malicious data into web traffic. Moreover, we conduct comprehensive studies involving 80 real-world Wi-Fi networks and 30 popular wireless routers. The results reveal that a majority of assessed Wi-Fi networks (75 out of 80) are vulnerable to our attack, and all tested routers fail to resist our attack. We have responsibly disclosed this vulnerability. While eliminating the side channel of encrypted frame size in Wi-Fi networks presents challenges, we propose several potential countermeasures to mitigate this vulnerability.

### v5

@@ V-B TCP Manipulation Attack
Experimental Setup. This attack involves three hosts: a web server, a victim client (our laptop), and an off-path attacker. We use a real financial website [^7] For ethical considerations, we anonymize this financial website in the paper. Moreover, our attack do not affect the website, since we only manipulate the web cache of the client side, i.e., our controlled laptop. as the web server. This website employs HTTP to deliver real-time Bitcoin price to the client in JSON format at 5-second intervals. Before launching the attack, the attacker uses the dig tool to probe the server’s IP address, explores the website to determine the server’s port, and familiarizes themselves with the JSON data structure. Consequently, the attacker can determine the server’s IP address and port, as well as identify specific data within the packet, as depicted in Figure 8(a), enabling manipulation of the victim’s web page. The victim client browses financial information on the web page via Wi-Fi. Both the attacker and the victim client are connected to the same Wi-Fi network. The off-path attacker attempts to detect and hijack the TCP connection between the victim client and the server. The server maintains a single long-lived TCP connection [^8] As recommended in RFC 2616 [40], the client typically does not maintain multiple long-lived TCP connections with the server simultaneously. with the client to transmit real-time financial data. Note that the modern browser may open multiple concurrent TCP connections along with the long TCP connection to speed up the page loading. These concurrent TCP connections are short-lived and have minimal influence on inferring the target long-lived TCP connection. Even if the server maintains multiple long-lived TCP connections with the client in some cases, the attacker can infer all the TCP connections and inject malicious data.
Attack Procedure. The web connection hijacking attack consists of five steps: (i) The attacker determines the MAC address and IP address of the victim client in the WLAN. (ii) By exploiting the encrypted frames, the attacker detects the client’s port number to obtain the TCP 4-tuple information. (iii) The attacker infers the exact sequence number and (iv) gets an acceptable acknowledgment number. (v) The attacker impersonates the server and injects forged TCP packets with the inferred values into the victim client. Finally, the client will accept the forged TCP packets, which subsequently update the financial information on the web page.
@@
Figure 9: Time/Bandwidth overheads of web manipulation.
Results Evaluation. Figure 9 illustrates the time cost and bandwidth consumption during the attack. It takes an average of 10.1 seconds to identify the client port number and 8.3 seconds to find the exact sequence number. Time cost required to find an acceptable acknowledgment number takes 9.6 seconds. The average duration of the entire attack is 28 seconds, with an average bandwidth cost of 46.32 KB/s. In this case, the attacker needs to infer an acceptable acknowledgment number, hence the success rate of this attack is lower than the TCP DoS attack but still exceeds 70%. After obtaining all the necessary information, the attacker sends forged TCP packets to the victim client and manipulates sensitive data on the web page. Figure 8(b) shows a snapshot of the manipulated web page and where attacker alters the Bitcoin price.
TABLE III: Details of 30 tested wireless routers.
@@
VI Real-world Attacks
To assess the impact of our attack, we conduct an extensive investigation on 30 popular wireless routers and 80 real-world Wi-Fi networks. We analyze 30 popular wireless routers and find that all the evaluated routers cannot protect the victim from our attack. Besides, we conduct SSH DoS attack and web hijack attack on the victim (i.e., our device) in the real-world Wi-Fi networks following the experimental setup and procedure in Section V. The results reveal that our attack is successful [^9] We conduct 10 iterations of the SSH DoS attack and the web hijack attack on the victim. In this context, the “successful” means that these two attacks can be successfully executed at least once in the real-world Wi-Fi network. against 75 (93.75%) of the 80 assessed Wi-Fi networks.
VI-A Analysis of AP Routers
@@
The early WEP used RC4 algorithm for data encryption, whereas WPA2 replaced them with the AES-CCMP algorithm. In the latest WPA3 standard, AES-GCMP is proposed to be used as the encryption method for WPA3 Enterprise mode. However, none of these security mechanisms can prevent the encrypted frame size from leaking upper layer information. We evaluate the wireless routers based on the experimental setup and attack procedures described in Section V. Based on our empirical findings, we confirm that all 30 evaluated wireless routers could not protect the supplicant from our attacks.
Table III shows detailed information on 30 tested wireless routers in our investigation. Take the first line for example, the evaluated router “Mi 4C” manufactured by Xiaomi belongs to the older Wi-Fi generation (Wi-Fi 4) and lacks support for IPv6 and WPA3. As outlined in the product description, the “Mi 4C” device offers support for various security features. These include a built-in firewall that allows administrators to define packet forwarding rules, a flood defense mechanism that restricts malicious flood traffic to prevent DoS attacks, and MAC address filtering, which enables network access authorization based on hardware addresses. In our investigation, all tested routers claim to support different security mechanisms to prevent various attacks. However, our study demonstrates that the existing security mechanisms are inadequate against our attack.
TABLE IV: Experimental results in 30 real-world Wi-Fi networks.
@@
VI-B Real-world Wi-Fi Networks Evaluation
The Wi-Fi scenarios we tested cover a wide range of public settings, including coffee shops, restaurants, hotels, cinemas, and bookstores. The experimental results illustrate that over 93% (i.e., 75 out of 80) of the evaluated Wi-Fi networks are vulnerable to our attack. By exploiting the encrypted frame size side-channel, the attacker can conduct SSH DoS and web hijacking attacks in the real-world Wi-Fi networks, achieving average success rates of 63.73% and 51.36% respectively in our evaluation. Next, we elaborate on the evaluation results.
@@ VI-B Real-world Wi-Fi Networks Evaluation
The 802.11n/ac standards are predominantly utilized (73.75%) in real-world Wi-Fi networks. This indicates that these Wi-Fi networks support two frequency bands (i.e., 2.4 GHz and 5 GHz), with their physical layer models based on the 802.11n and 802.11ac standards, respectively. There is only 17.5% (14 out of 80) of the evaluated Wi-Fi networks support 802.11ax. This is consistent with our expectations, as the 802.11ax standard was defined in 2019 and would require more time for widespread deployment. Furthermore, the simultaneous utilization of two channels in Wi-Fi networks is the most prevalent case (76.25%) due to the widespread support and default configuration of dual-channel capabilities in wireless routers. In certain scenarios, such as office buildings, Wi-Fi networks employ multiple wireless channels to enhance network performance. In our study, we identify 12 Wi-Fi networks that utilize multiple wireless channels. At first glance, the usage of multiple wireless channels might appear as a minor hurdle to our attack, as the attacker needs to perform additional channel scanning to determine the specific channel employed by the victim. Indeed, the attacker can enhance the success rate by employing a channel “eviction” strategy, which will be explained in detail in Section VII. Additionally, we encounter several Wi-Fi networks that operate on a single channel. When questioned, network administrators cited security considerations as the rationale behind this choice, although they did not provide any further specifics.
@@
Out of all the evaluated networks, 16 (20%) of them are open and do not encrypt the supplicant’s data frames. This means that an attacker can potentially access and view the contents of all supplicant’s frames transmitted on these networks, representing a significant breach of supplicant privacy. The remaining 64 (80%) Wi-Fi networks utilize WPA2/WPA3 to encrypt the supplicant’s wireless frame. It is worth noting that out of the 80 Wi-Fi networks evaluated, 11 (13.75%) of them have AP isolation enabled. In these Wi-Fi networks, the attacker obtains the victim’s MAC and IP address by leveraging the DHCP mechanism and injects malicious packets into the victim through the AP’s external port [^11] We do not encounter Wi-Fi networks with both AP isolation enabled and MFP required in our evaluations.. However, our attack encounters failure in five Wi-Fi networks. Among them, one network is equipped with reverse path authentication [55, 62], preventing the attacker from sending packets from the WLAN to the AP’s external port. In the remaining four Wi-Fi networks, the attacker cannot obtain the AP’s external port as the gateway does not respond to ICMP ping messages.
@@
We elaborate on the experimental results of 30 encrypted Wi-Fi networks in Table IV. Details of all 80 evaluated Wi-Fi networks are presented in Appendix A. We take the first row of Table IV as an example to analyze the results. In our study, the SSID “Bookstore 1” indicates a Wi-Fi network that is accessible in a bookstore. It is common practice to set the Wi-Fi SSID as the organization name, which may expose the organization’s identity. Therefore, to protect anonymity, we have anonymized the Wi-Fi SSID in this paper. This bookstore’s Wi-Fi network only supports IPv4 and does not have AP isolation enabled, while its AP is produced by ADSLR. This AP provides two access channels (i.e., 6 and 161) and employs the 802.11n and 802.11ac standards. The TCP connection of the victim supplicant can be hijacked using the attack presented in Section IV. The success rates for conducting SSH DoS and web hijacking on the victim supplicant are 70% and 60%, respectively.
@@
Within the evaluated vulnerable Wi-Fi networks, we have observed a range of success rates for our attack, varying from 30% to 80%. The principal factor influencing this variance is the heterogeneous wireless environments (e.g., different wireless interference and channel contention) encountered in real-world Wi-Fi networks, leading to varying capabilities for attacker to capture the victim’s Wi-Fi frames. Factors such as wireless interference (e.g., from microwave ovens and Bluetooth devices) and channel contention can hinder the attacker’s ability to capture the victim’s Wi-Fi frames, resulting in the failure of attacks. For instance, as illustrated in Table IV, Coffee Shop 1, situated on a university campus, experiences less wireless interference and channel contention compared to Restaurant 3, located within a large shopping mall. Consequently, the success rate of attacks in the latter Wi-Fi network is lower due to the elevated interference and contention in that environment. We will conduct a more comprehensive analysis of the factors influencing the success of our attack in Section VII.
VII Discussion
@@
Channel Contention. APs and supplicants based on the 802.11 standards use Carrier Sense Multiple Access with Collision Avoidance (CSMA/CA) to compete equally for the occupation of the wireless channel. Before transmitting frames, wireless channel listening is conducted to ensure that the channel is not occupied. Frames are transmitted only after verifying the channel’s availability. Due to channel contention, there may be an uncertain delay or even frame dropping in the victim’s responses to the probe packets. This uncertain response delay or frame dropping is the primary reason for the fluctuating success rate of our attack because the attacker needs to analyze the victim’s encrypted frames within a time slice after the probe packets are sent. To mitigate channel contention, we propose a channel “eviction” strategy. The attacker can evict other supplicants from the channel used by the victim. Specifically, the attacker impersonates the AP and sends decertification frames to the supplicant, causing it to detach from the current channel of the AP [^12] In the Wi-Fi network with MFP (Management Frame Protection) enabled, attackers can exploit implementation vulnerabilities to force the supplicant to detach from the current channel [53, 52].. The supplicant will attempt to reconnect to the Wi-Fi network, but after encountering several disconnections, it will switch to another channel. This strategy requires the Wi-Fi network to support multiple access channels. Fortunately, most Wi-Fi networks provide more than one access channel, as shown in Section VI-B. Note that the channel switching (i.e., our “eviction” strategy to cause other supplicants to detach from the current channel) is transparent to the users. The only impact is that the user may experience a brief (a few seconds) network jitter during the channel switching.
@@ IX Related Work
In TCP connection hijacking attacks, side channels serve as potent tools for attackers. The IPID, in particular, has been a frequent target for exploitation. For example, Jeffrey et al. utilized per-destination IPID counters to estimate the number of packets transmitted between two machines and even detect the presence of a TCP connection [36]. Similarly, Alexander et al. used the IPID of triggered RST packets to identify the existence of the victim’s TCP connection [8]. In a recent instance, Feng et al. manipulated the IPID assignment using ICMP to hijack the victim’s TCP connection [23]. Besides IPID, the challenge ACK mechanism is another side channel exploited by off-path attackers. Cao et al., for example, utilized the global rate limit of challenge ACK to infer and hijack TCP connections [15, 16]. Moreover, a timing side channel has been found in half-duplex Wi-Fi technology, which can be exploited by off-path attackers to inject data into the victim’s TCP connection [17]. However, this method typically requires attackers to install a puppet on the victim client, which is not necessary for our attack. Tolley et al. recently proposed a blind in/on-path attack in VPNs, aiming to infer the existence of, interfere with, or inject data into TCP connections forwarded through encrypted VPN tunnels [65]. Different from previous research, our work reveals a new side channel, i.e., information leakage due to the encrypted frame size in Wi-Fi networks. This side channel can be exploited by a pure off-path attacker to hijack victim’s TCP connections.
X Conclusion
@@
In this paper, we present a new off-path TCP hijacking attack that takes advantage of the encrypted frame size in Wi-Fi networks to detect and hijack TCP connections belonging to a victim supplicant. Our attack focuses on TCP connection hijacking, but non-TCP sessions may also be affected due to Wi-Fi layer information leakage, which we will explore in future work. This side channel (i.e., the observable frame size) vulnerability is an inherent flaw in Wi-Fi networks, specifically the 802.11 standards. To execute our attack, attackers initially scan the WLAN to identify active victim supplicant, then analyze the victim’s encrypted frame size to infer the 4-tuple, exact sequence number, and acceptable acknowledgment number of the victim’s TCP connection. Specifically, our attacker can hijack the victim’s TCP connection within 28 seconds. We carry out our attack in typical Wi-Fi scenarios, and our evaluation demonstrates that this new off-path TCP hijacking attack can result in significant damage to upper-layer applications, such as SSH DoS and the injection of malicious data into web traffic. Moreover, we conduct comprehensive studies involving 80 real-world Wi-Fi networks and 30 popular wireless routers. The results reveal that a majority of assessed Wi-Fi networks (75 out of 80) are vulnerable to our attack, and all tested routers fail to resist our attack. We have responsibly disclosed this vulnerability. While eliminating the side channel of encrypted frame size in Wi-Fi networks presents challenges, we propose several potential countermeasures to mitigate this vulnerability.
Acknowledgment
We thank the anonymous reviewers for their thoughtful comments. This work was in part supported by NSFC under U22B2025, Jiangsu Provincial Scientific Research Center of Applied Mathematics under Grant No. BK20233002, the National Science Foundation for Distinguished Young Scholars of China under No. 62425201, the Science Fund for Creative Research Groups of the National Natural Science Foundation of China under No. 62221003, the Key Program of the National Natural Science Foundation of China under No. 61932016 and No. 62132011. Ziqiang Wang, Ke Xu and Jianping Wu are corresponding authors.
Appendix A The 80 evaluated Wi-Fi networks
All the 80 tested Wi-Fi networks are shown in Table V.
No.
SSID
AP Vendor
IPv4/IPv6
PHY model
AP isolation
W
i-Fi channel
SSH DoS
Web hijack
1 Bookstore 1 ADSLR 802.11n/ac No 6, 161 7/10 6/10
2 Bookstore 2 HUAWEI 802.11n/ac/ax No 11, 44 7/10 7/10
[…]
4 Bookstore 4 Xiaomi 802.11n/ac No 6, 149 8/10 7/10
5 Bookstore 5 TP-LINK 802.11n/ac No 7, 36 7/10 5/10
6 Bookstore 6 Tenda 802.11n/ac No 9, 48 5/10 3/10
7 Bookstore 7 Ruijie 802.11n/ac No 5, 149 6/10 5/10
8 Coffee Shop 1 TP-LINK 802.11n/ac No 6, 60 8/10 6/10
9 Coffee Shop 2 HUAWEI 802.11n/ac No 1, 36 8/10 7/10
10 Coffee Shop 3 Wimaster 802.11n/ac Yes 1, 48 7/10 6/10
11 Coffee Shop 4 Tenda 802.11n/ac No 4, 153 6/10 5/10
12 Coffee Shop 5 TP-LINK 802.11n/ac No 153 7/10 5/10
[…]
14 Coffee Shop 7 Xiaomi 802.11n/ac No 8, 36 6/10 6/10
[…]
16 Restaurant 1 D-Link 802.11n/ac No 5, 149 7/10 5/10
17 Restaurant 2 TP-LINK 802.11n/ac No 1, 153 6/10 6/10
18 Restaurant 3 Ruijie 802.11n/ac Yes 11, 64 6/10 4/10
19 Restaurant 4 iKuai 802.11n/ac No 1, 48 5/10 3/10
[…]
21 Restaurant 6 Xiaomi 802.11n/ac No 36 6/10 5/10
[…]
23 Restaurant 8 HUAWEI 802.11n/ac No 11, 157 5/10 4/10
24 Office building 1 TP-LINK 802.11n/ac No 11, 36, 40 7/10 6/10
25 Office building 2 H3C 802.11n/ac No 1, 48, 153 8/10 7/10
26 Office building 3 Netcore 802.11n/ac Yes 6, 149 8/10 6/10
27 Office building 4 ZTE 802.11n/ac No 11, 60 6/10 6/10
[…]
29 Office building 6 Linksys 802.11n/ac/ax No 11, 48 6/10 4/10
30 Office building 7 HUAWEI 802.11n/ac No 9, 161 7/10 5/10
31 Office building 8 TP-LINK 802.11n/ac No 1, 157 7/10 6/10
32 Enterprise 1 TP-LINK 802.11n/ac No 6, 36 6/10 6/10
33 Enterprise 2 HUAWEI 802.11n/ac Yes 11, 157 7/10 6/10
34 Enterprise 3 Ruijie 802.11n/ac Yes 1, 11, 40, 149 6/10 5/10
[…]
37 Enterprise 6 PHICOMM 802.11n/ac No 6, 36 6/10 4/10
[…]
40 Fast Food Restaurant 2 TP-LINK 802.11n/ac No 3, 157 7/10 6/10
[…]
42 Fast Food Restaurant 4 Ruijie 802.11n/ac No 1, 44 6/10 6/10
43 Fast Food Restaurant 5 HUAWEI 802.11n/ac/ax No 6, 153 5/10 4/10
[…]
45 Fast Food Restaurant 7 Xiaomi 802.11n/ac No 1, 52 6/10 5/10
46 Fast Food Restaurant 8 ZTE 802.11n/ac No 11, 40 5/10 3/10
47 Cinema 1 HUAWEI 802.11n/ac No 1, 157 7/10 6/10
[…]
49 Cinema 3 Ruijie 802.11n No 6 7/10 6/10
50 Cinema 4 H3C 802.11n/ac No 10, 149 7/10 5/10
[…]
52 Hotel 1 HUAWEI 802.11n/ac No 6, 44 8/10 7/10
53 Hotel 2 Ruijie 802.11n No 1, 11 8/10 6/10
54 Hotel 3 D-Link 802.11n/ac No 1, 48 6/10 5/10
55 Hotel 4 Xiaomi 802.11n Yes 1 5/10 4/10
56 Hotel 5 TP-LINK 802.11n/ac No 9, 48 6/10 5/10
[…]
58 Hotel 7 HUAWEI 802.11n/ac No 60 6/10 4/10
59 Hotel 8 Wimaster 802.11n/ac/ax No 6, 56 7/10 5/10
60 Experience Store 1 HUAWEI 802.11n/ac No 1, 36 7/10 6/10
61 Experience Store 2 HUAWEI 802.11n/ac No 11, 149 7/10 6/10
62 Experience Store 3 Tenda 802.11n/ac No 4,153 6/10 5/10
63 Experience Store 4 TP-LINK 802.11n/ac No 11, 36 5/10 3/10
[…]
65 Experience Store 6 H3C 802.11n/ac/ax No 8, 52 5/10 4/10
[…]
67 Campus 1 Xiaomi 802.11n/ac No 9, 36 6/10 4/10
68 Campus 2 Ruijie 802.11n/ac No 1, 44 7/10 6/10
69 Campus 3 H3C 802.11n/ac No 1, 6, 40, 64 6/10 6/10
70 Campus 4 ASUS 802.11n/ac/ax No 6, 40 5/10 3/10
[…]
72 Campus 6 Netgear 802.11n/ac/ax No 3, 149 6/10 5/10
[…]
74 Shopping Mall 1 Ruijie 802.11n/ac No 11, 149 4/10 3/10
[…]
76 Shopping Mall 3 HUAWEI 802.11n/ac No 1, 157 6/10 6/10
[…]
78 Shopping Mall 5 HUAWEI 802.11n/ac No 1, 11, 48, 157 7/10 5/10
79 Shopping Mall 6 H3C 802.11n/ac/ax No 1, 36, 44 6/10 4/10
[…]
  • ◐ means IPv4 only and ● means both IPv4 and IPv6 are supported.
[…]
    1F5F8 indicates the WiFi network does not encrypt frames, letting the attacker obtain the victim’s TCP connection information directly.
[…]
    2717 indicates the attack failed in the Wi-Fi network.
TABLE V: Experimental results in 80 real-world Wi-Fi networks.